Allison:

Evaluate d - (-3.5) when d = 2.74

Layla:

your answer is -0.76

Allison:

Oh I got positive .76 thank youuuu

Layla:

welcome

Vocaloid:

Replace d with 2.74 2.74 - (-3.5) = ?

Layla:

because 2.74 minus -3.5 = -0.76

Allison:

Thank you c: @Vocaloid and @Layla

Vocaloid:

2.74 - (-3.5) = 2.74 + 3.5 = ?

Vocaloid:

Two negative signs together means +

Layla:

yes but there is only 1 negetive sign so it will be a negetive

Vocaloid:

Use a calculator if you don't believe me.

Layla:

no, it says evaluate D MINUS negitive 3.5

Vocaloid:

1 attachment
Layla:

did you seriuosly just look up the answer?!? The rules say to use your brain

Layla:

I I don't mean to be mean or anything

Allison:

Layla calm down she's just helping prove a point.

Layla:

I just said that I wasn't trying to be mean I just asked if she looked it up

Allison:

It's 6.2 right?

Vocaloid:

Almost

Allison:

They're both negative numbers omg.

Layla:

6.24

Layla:

you were only 0.04 points off

Allison:

http://prntscr.com/gn714t

Layla:

yes so you are right

Vocaloid:

Ththats not the same problem

Vocaloid:

That's

Allison:

Did I type it wrong?

Layla:

probably XD

Allison:

Hold up.

Allison:

*Kms*

Allison:

*Kms*

Allison:

I'm looking at a completely different problem. *FACEPALM*

Layla:

omg lol XD